### v5

# {cutout}00.3cm22cm1 empty    PointMamba: A Simple State Space Model for Point Cloud Analysis

Abstract
Transformers have become one of the foundational architectures in point cloud analysis tasks due to their excellent global modeling ability. However, the attention mechanism has quadratic complexity, making the design of a linear complexity method with global modeling appealing. In this paper, we propose PointMamba [Uncaptioned image], transferring the success of Mamba, a recent representative state space model (SSM), from NLP to point cloud analysis tasks. Unlike traditional Transformers, PointMamba employs a linear complexity algorithm, presenting global modeling capacity while significantly reducing computational costs. Specifically, our method leverages space-filling curves for effective point tokenization and adopts an extremely simple, non-hierarchical Mamba encoder as the backbone. Comprehensive evaluations demonstrate that PointMamba achieves superior performance across multiple datasets while significantly reducing GPU memory usage and FLOPs. This work underscores the potential of SSMs in 3D vision-related tasks and presents a simple yet effective Mamba-based baseline for future research. The code will be made available at https://github.com/LMD0311/PointMamba.
Figure 1: Comprehensive comparisons between our PointMamba and its Transformer-based counterparts liu2022masked ; chen2023pointgpt ; dong2022autoencoders . (a) Without bells and whistles, our PointMamba achieve better performance than the representative Transformer-based methods on the various point cloud analysis datasets. (b)-(d) The Transformer presents quadratic complexity, while our PointMamba has linear complexity. For example, with the length of point tokens increasing, we significantly reduce GPU memory usage and FLOPs and have a faster inference speed compared to the most convincing Transformer-based method, i.e., PointMAE liu2022masked .
1 Introduction
Point cloud analysis is one of the fundamental tasks in computer vision and has a wide range of real-world applications shi2020points ; wu2023point ; chen2023voxelnext , including robotics, autonomous driving, and augmented reality. It is a challenging task due to the intrinsic irregularity and sparsity of point clouds. To address the issues, there has been rapid progress in deep learning-based methods pang2022masked ; liu2022masked ; qi2017pointnet++ ; qian2022pointnext , consistently pushing the performance to the new record.
Recently, Transformers have achieved remarkable progress in point cloud analysis. The key to the Transformer is the attention mechanism, which can effectively capture the relationship of a set of points. By integrating self-supervised learning paradigms with fine-tuning for downstream tasks, these Transformer-based methods have achieved superior performance yu2022point ; liu2022masked ; qi2023recon . However, the complexity of attention mechanisms is quadratic, bringing significant computational cost, which is not friendly to low-resource devices. Thus, this naturally raises a question: how to design a simple, elegant method that operates with linear complexity, thereby retaining the benefits of global modeling for point cloud analysis?
We note the recent advance of the State Space Models (SSMs). As a pioneer, the Structured State Space Sequence Model gu2021efficiently (S4) has emerged as a promising class of architectures for sequence modeling thanks to its strong representation ability and linear-time complexity (achieved by eliminating the need to store the complete context). Another pioneer, Mamba gu2023mamba , adopts time-varying parameters to the SSM based on S4, proposing an efficient hardware-aware algorithm to enable highly efficient training and inference with dynamic modeling. Recently, a few concurrent methods zhu2024vision ; liu2024vmamba successfully transfer the 1D-sequence Mamba from NLP to 2D vision tasks (e.g., image classification and segmentation), achieving similar or surpass the Transformer counterpart dosovitskiy2020image while significantly reducing memory usage. However, regarding the more complex, unstructured data, e.g., 3D point cloud, the effectiveness of Mamba remains unclear. The lack of early exploration of Mamba’s potential for point cloud-related tasks hinders further development of its capabilities across the diverse range of applications in this domain.
Inspired by this, this paper aims to unlock the potential of SSM in point cloud analysis tasks, discussing whether it can be a viable alternative to Transformers in this domain. Through a series of pilot experiments, we find that directly using the pioneering SSM, Mamba gu2023mamba , can not achieve ideal performance. We argue that the main inherent limitation comes from the unidirectional modeling employed by the default Mamba, as the context is obtained by compressing the historical hidden state instead of through the interaction between each element. In contrast, the self-attention of the Transformer is invariant to the permutation of the input elements. Given the three-dimensional nature (e.g., unstructured and disordered) of point clouds, using a single scanning process often struggles to concurrently capture dependency information across various directions, which makes it difficult to construct global modeling for the RNN-like modes (e.g., Mamba).
Therefore, we introduce a simple yet effective Point Cloud State Space Model (denoted as PointMamba [Uncaptioned image]) with global modeling and linear complexity. Specifically, to enable Mamba to capture the point cloud structure causally, we first use a point tokenizer to generate two types of point tokens via a point scanning strategy, employing two space-filling curves to scan key points from different directions. As a result, the unstructured 3D point clouds can be transformed into a regular sequence. The first type of token has local modeling capabilities through sequential encoding, with the latest token holding global sequence information. Consequently, the second type of token can achieve global modeling by containing global information that comes from the first type. Besides, we propose an extremely simple order indicator to maintain the distinct spatial characteristics of different scanning when training, preserving the integrity of the spatial information.
To make the model as simple as possible, PointMamba only employs plain and non-hierarchical Mamba as the backbone to extract features for given serialized point tokens without bells and whistles. We demonstrate that PointMamba is very flexible in the pre-training paradigm, where we customize an MAE-like pertaining strategy to provide a good prior, which chooses a random serialization strategy from a pre-defined serialization bank to perform mask modeling. It facilitates the model to exact the general local relationships from different scanning perspectives, better matching the requirement of indirection modeling of mamba.
Despite no elaborate or complex designs in the model, our PointMamba achieves superior performance on various point cloud analysis datasets (Fig. 1(a)). Besides the superior performance, thanks to the linear complexity of Mamba, we show the surprisingly low computational cost [^1] Note that we removed the tokenizer of both Point-MAE and PointMamba , directly fed with a predefined sequence, to better illustrate the structural efficiency., as shown in Fig. 1(b)-(c). These notable results underscore the potential of SSM in 3D vision-related tasks.
In conclusion, the contributions of this paper are twofold. 1) We introduce the first state space model for point cloud analysis, named PointMamba [Uncaptioned image], which features global modeling with linear complexity. Despite the absence of elaborate or complex structural designs, PointMamba demonstrates its potential as an optional model for 3D vision applications. 2) Our PointMamba exhibits impressive capabilities, including structural simplicity (e.g., vanilla Mamba), low computational cost, and knowledge transferability (e.g., support for self-supervised learning).
2 Related work
2.1 Point Cloud Transformers
Vision Transformer (ViT) dosovitskiy2020image has become one of the mainstream architectures in point cloud analysis tasks due to its excellent global modeling ability. Specifically, Point-BERT yu2022point and Point-MAE pang2022masked introduce a standard Transformer architecture for self-supervised learning and is applicable to different downstream tasks. Several works further introduce GPT scheme chen2023pointgpt , multi-scale zhang2022point ; zha2024towards , and multi-modal dong2022autoencoders ; qi2023recon ; qi2024shapellm to guide 3D representation learning. On the other hand, some researchers guo2021pct ; zhao2021point ; wang2023octformer focus on modifying the Transformers for point clouds. The PCT guo2021pct conducts global attention directly on the point cloud. Point Transformer zhao2021point applies vector attention zhao2020exploring to perform local attention between each point and its adjacent points. The later Point Transformer series wu2022point ; wu2023point further extends the performance and efficiency of the Transformer for different tasks. OctFormer wang2023octformer leverages sorted shuffled keys of octrees to partition point clouds and significantly improve efficiency and effectiveness.
Standard Transformers can be smoothly integrated into autoencoders using an encoder-decoder design he2022masked , which makes this structure ideal for pre-training and leads to significant performance improvements in downstream point cloud analysis tasks yu2022point ; pang2022masked ; chen2023pointgpt ; zha2024lcm ; zha2023instance . However, the attention mechanism has a time complexity of $O(n^{2}d)$, where $n$ represents the length of the input token sequence and $d$ represents the dimension of the Transformer. This implies that as the input sequence grows, the operational efficiency of the Transformer is significantly constrained.
In this work, we focus on designing a simple State Space Model (SSM) for point cloud analysis without attention while maintaining the global modeling advantages of the Transformer.
2.2 State Space Models
Linear state space equations gu2020hippo ; gu2021combining , combined with deep learning, offer a compelling approach for modeling sequential data, presenting an alternative to CNNs or Transformers. The Structured State Space Sequence Model gu2021efficiently (S4) leverages a linear state space for contextualization and shows strong performance on various sequence modeling tasks, especially with lengthy sequences. To alleviate computational burden, HTTYH gu2022train , DSS gupta2022diagonal , and S4D gu2022parameterization propose employing a diagonal matrix within S4, maintaining performance without excessive computational costs. The S5 smith2022simplified proposes a parallel scan and the MIMO SSM, enabling the state space model to be efficiently utilized and widely implemented. Recently, Mamba gu2023mamba introduced the selective SSM mechanism, a breakthrough achieving linear-time inference and effective training using a hardware-aware algorithm, garnering considerable attention. In the vision domain, Vision Mamba liu2024vmamba compresses the visual representation through bidirectional state space models. VMamba zhu2024vision introduces the Cross-Scan Module, enabling 1D selective scanning in 2D images with global receptive fields. Besides, the great potential of Mamba motivates a series of work across diverse domains, including graph wang2024graph ; behrouz2024graph , medical segmentation ma2024u ; liu2024swin , video understanding li2024videomamba ; chen2024video and generative models hu2024zigma ; li2024mamba .
To the best of our knowledge, there are limited works that study SSMs for point cloud analysis. In this work, we delve into the potential of Mamba in point cloud analysis and propose PointMamba, which achieves superior performance and significantly reduces computational costs.
3 Preliminaries
State Space Model. Drawing inspiration from control theory, the State Space Model (SSM) represents a continuous system that maps a state $x_{t}$ to $y_{t}$ through an implicit latent state $h_{t}\in\mathbb{R}^{N}$. To integrate SSMs into deep models, S4 gu2021efficiently defines the system with four parameters ($\bm{A},\bm{B},\bm{C},$ and sampling step size $\bm{\Delta}$). The sequence-to-sequence transformation is defined as:
$h_{t}=\overline{\bm{A}}h_{t-1}+\overline{\bm{B}}x_{t},\qquad y_{t}=\bm{C}h_{t}+\bm{D}x_{t},$ (1)
where $\bm{C}\in\mathbb{R}^{1\times N}$ is a project parameter, and $\bm{D}\in\mathbb{R}^{1\times N}$ represents a residual connection. The parameters $\overline{\bm{A}},\overline{\bm{B}}$ are defined using the zero-order hold (ZOH) discretization rule:
$\overline{\bm{A}}\in\mathbb{R}^{N\times N}=\exp(\bm{A\Delta}),\qquad\overline{\bm{B}}\in\mathbb{R}^{N\times 1}=\left(\bm{A\Delta}\right)^{-1}\left(\exp(\bm{A\Delta})-\bm{I}\right)\cdot\bm{\Delta B}.$ (2)
However, the parameter ($\overline{\bm{A}},\overline{\bm{B}},\bm{C},\bm{\Delta}$) are fixed across all time steps due to the Linear Time-Invariant (LTI) property of SSMs, which limits their capacity to handle varied input sequences.
Recently, Selective SSM (S6) considers parameters $\bm{B},\bm{C},\bm{\Delta}$ as functions of the input, effectively transforming the SSM into a time-variant model. Our PointMamba adopts a hardware-aware implementation gu2023mamba of S6, showing linear complexity and strong sequence modeling capability.
Space-filling curve. Space-filling curves are paths that traverse every point within a higher-dimensional discrete space while maintaining spatial proximity to a certain degree. Mathematically, they can be defined as a bijective function $\Phi:\mathbb{Z}\to\mathbb{Z}^{3}$ for point clouds. Our PointMamba focuses on the Hilbert space-filling curve hilbert1935stetige and its transposed variant (called Trans-Hilbert), both of which are recognized for effectively preserving locality, ensuring that data points close in $\mathbb{Z}$ space remain close after transformation to $\mathbb{Z}^{3}$. We note that some methods wu2023point ; wang2023octformer utilize space-filling curves to partition the point cloud for capturing spatial contexts, whereas our work mainly focuses on transferring the point clouds to serialization-based sequences and combine with Mamba to implement global modeling. The motivation and objective are different.
4 PointMamba
This paper aims to design a simple yet solid Mamba-based Point cloud analysis method. The pipeline of our method is shown in Fig. 2. Starting with an input point cloud, we first sample the key points via Farthest Point Sampling (FPS). Then, a simple space-scanning strategy is applied to reorganize these points, resulting in serialized key points. Under a KNN and lightweight PointNet qi2017pointnet , we obtain the serialized point tokens. Finally, the entire sequence is subsequently processed by a plain, non-hierarchical encoder structure composed of several stacked Mamba blocks. Besides, to provide a good prior for PointMamba, we propose a serialization-based mask modeling paradigm, which randomly chooses a space-filling curve for serialization and mask, as shown in Fig. 4.
Figure 2: The pipeline of our PointMamba. It is simple and elegant, without bells and whistles. We first utilize Farthest Point Sampling (FPS) to select the key points. Then, we propose to utilize two types of space-filling curves, including Hilbert and Trans-Hilbert, to generate the serialized key points. Based on these, the KNN is used to form point patches, which will be fed to the token embedding layer to generate the serialized point tokens. To indicate the tokens generated from which space-filling curve, the order indicator is proposed. The encoder is extremely simple, consisting of $N\times$ plain and non-hierarchical Mamba blocks.
4.1 The structure of PointMamba
In this section, we introduce the structure of our PointMamba. The goal of this paper is to provide a simple yet solid Mamba baseline for point cloud analysis tasks and explore the potential of plain and non-hierarchical Mamba. Thus, in the spirit of Occam’s razor, we make the structure as simple as possible without any complex or elaborate design.
Point scanning strategy. Building on the pioneer works yu2022point ; pang2022masked , we first utilize the Farthest Point Sampling (FPS) to select the key points. Specifically, given an input point cloud $\bm{P}\in\mathbb{R}^{M\times 3}$, where $M$ is the number of points, the FPS is applied to sample $n$ key points from the original point cloud $\bm{P}$, resulting in $\bm{p}\in\mathbb{R}^{n\times 3}$. In general, the order of the sampling key points $\bm{p}$ is random, without specific order. This is not a significant problem for the previous Transformer-based methods, as the Transformer is order-invariant when processing sequence data: in the self-attention mechanism, each element at a given position can interact with all other elements in the sequence through attention weights. However, for the selective state space model, i.e., Mamba, we argue that it is hard to model the unstructured point clouds due to the unidirectional modeling. Thus, we propose to leverage the space-filling curves to transform the unstructured point clouds into a regular sequence. Specifically, we choose two representative space-filling curves to scan the key points: the Hilbert curve hilbert1935stetige and its transposed variant, denoted as Trans-Hilbert. Compared with the random sequence, space-filling curves like the Hilbert curve can preserve spatial locality well, i.e., along the scanned 1D serialized point sequence, adjacent key points often have geometrically close positions in 3D space. We argue that this property ensures that the spatial relationships between points are largely maintained, which is crucial for accurate feature representation and analysis in point cloud data. As a complementary, the Trans-Hilbert performs similarly but scans from different clues, which can provide diverse perspectives on spatial locality. By applying Hilbert and Trans-Hilbert to the key points, we obtain two different point serializations, $\bm{p}_{h}$ and $\bm{p}_{h^{\prime}}$, which will be used to construct point tokens.
Point tokenizer . After obtaining the two serialized key points $\bm{p}_{h}$ and $\bm{p}_{h^{\prime}}$, we then utilize the KNN algorithm to select $k$ nearest neighbors for each key point, forming $n$ token patches $\bm{T}_{h}\in\mathbb{R}^{n\times k\times 3}$ and $\bm{T}_{h^{\prime}}\in\mathbb{R}^{n\times k\times 3}$ with patch size $k$. To aggregate local information, points within each patch are normalized by subtracting the key point to obtain relative coordinates. We map the unbiased local patches to feature space using a lightweight PointNet qi2017pointnet (point embedding layer), obtaining serialized point tokens $\bm{E}^{h}_{0}\in\mathbb{R}^{n\times C}$ and $\bm{E}^{h^{\prime}}_{0}\in\mathbb{R}^{n\times C}$, where the former is the Hilbert-based and the latter is Trans-Hilbert-based.
Order indicator. Directly fed the two type serialized point tokens $\bm{E}^{h}_{0}\in\mathbb{R}^{n\times C}$ and $\bm{E}^{h^{\prime}}_{0}\in\mathbb{R}^{n\times C}$ into Mamba encoder might cause confusion as $\bm{E}^{h}_{0}$ and $\bm{E}^{h^{\prime}}_{0}$ actually share the same center but with different order. Maintaining the distinct characteristics of these different scanning strategies is important for preserving the integrity of the spatial information. Thus, we propose an extremely simple order indicator to indicate the scanning strategy used. Specifically, the proposed order indicator performs the linear transformation to transfer features into different latent spaces. The formulation can be written as follows:
$\bm{Z}^{h}_{0}=\bm{E}^{h}_{0}\odot\bm{\gamma}_{h}+\bm{\beta}_{h},\quad\bm{Z}^{h^{\prime}}_{0}=\bm{E}^{h^{\prime}}_{0}\odot\bm{\gamma}_{h^{\prime}}+\bm{\beta}_{h^{\prime}},$ (3)
where $\bm{\gamma}_{h}$/$\bm{\gamma}_{h^{\prime}}\in\mathbb{R}^{C}$ and $\bm{\beta}_{h}$/$\bm{\beta}_{h^{\prime}}\in\mathbb{R}^{C}$ refer to the scale and shift factors, respectively. $\odot$ is the dot product. We then concat $\bm{Z}^{h}_{0}$ and $\bm{Z}^{h^{\prime}}_{0}$, resulting in $\bm{Z}_{0}\in\mathbb{R}^{2n\times C}$.
Mamba encoder. After obtaining the token $\bm{Z}_{0}$, we will feed it into the encoder, containing N $\times$ Mamba block, to extract the feature. Specifically, for each Mamba block, layer normalization ($\mathrm{LN}$), Selective SSM, depth-wise convolution chollet2017xception ($\mathrm{DW}$), and residual connections are employed. A standard Mamba layer is shown in Fig. 2, and the output can be summarized as follows:
$\bm{Z}^{\prime}_{l-1}=\mathrm{LN}\left(\bm{Z}_{l-1}\right),\qquad\quad\;\bm{Z}^{\prime}_{l}=\sigma\left(\mathrm{DW}\left(\mathrm{Linear}\left(\bm{Z}^{\prime}_{l-1}\right)\right)\right)$ (4)
$\bm{Z}^{\prime\prime}_{l}=\sigma\left(\mathrm{Linear}\left(\bm{Z}^{\prime}_{l-1}\right)\right),~{}~{}\enspace\bm{Z}_{l}=\mathrm{Linear}\left(\mathrm{SelectiveSSM}\left(\bm{Z}^{\prime}_{l}\right)\odot\bm{Z}^{\prime\prime}_{l}\right)+\bm{Z}_{l-1}$
Figure 3: An intuitive illustration of global modeling from PointMamba.
$\bm{Z}_{l}\in\mathbb{R}^{2n\times C}$ is the output of the $l$-th block, and $\sigma$ indicates $\mathrm{SiLU}$ activation hendrycks2016gaussian . The $\mathrm{SelectiveSSM}$ is the key to the Mamba block, with a detailed description in Sec. 3. To better understand why the proposed PointMamba has global modeling capacity, we provide an intuitive visualization. As shown in Fig. 3, after modeling the first group of point tokens (i.e., Hilbert-based), the accumulated global information can improve the serialization process for the next set of tokens (i.e., Trans-Hilbert-based). This mechanism ensures that each serialized point in the Trans-Hilbert sequence is informed by the entire history of the previously processed Hilbert sequence, thereby enabling a more contextually rich and globally aware modeling process. More discussions can be found in Appendix A.2.
In our study, we show that even a very simple Mamba block without specific designs, our PointMamba can surpass the various Transformer-based point cloud analysis methods.
4.2 The serialization-based mask modeling
Figure 4: The details of our proposed serialization-based mask modeling. During the pre-training, we randomly choose one space-filling curve to generate the serialized point tokens for mask modeling, and different serialized point tokens have different order indicators.
One intriguing characteristic of Transformers-based methods liu2022masked ; yu2022point ; chen2023pointgpt is their improved performance using the pre-training scheme, especially mask modeling he2022masked . In this paper, considering the unidirectional modeling of Mamba, we customize a simple yet effective serialization-based mask modeling paradigm, as shown in Fig. 4.
Specifically, after obtaining the key points, we randomly choose Hilbert or Trans-Hilbert curve to implement serialization in each iteration, resulting in serialization-based key points, i.e., $\bm{p}_{h}$ and $\bm{p}_{h^{\prime}}$, are obtained from Hilbert and Trans-Hilbert, respectively. Such a scheme allows the model to exact the local relationships from different scanning clues. Then, the KNN and the token embedding layer are used to generate the point tokens. To discriminate the point tokens serialized from which space-filling curves, we apply the order indicator to the point tokens, where different serialized point tokens have different order indicators, which are similar to the mentioned Eq. 3. Next, we randomly mask the serialization-based point tokens with a high ratio of 60%. Then, an asymmetric autoencoder, consisting of several vanilla Mamba blocks, is employed to extract the point feature, and the final layer of the autoencoder utilizes a simple prediction head for reconstruction. To reconstruct masked point patches in coordinate space, we employ a linear head to project the masked token to the shape of the masked input points. The Chamfer Distance fan2017point is then used as the reconstruction loss to recover the coordinates of the points in each masked point patch.
We demonstrate that with such a simple serialization-based mask modeling paradigm, PointMamba can easily achieve superior performance.
5 Experiments
5.1 Implementation details
SSM is new to 3D point cloud analysis, with no existing works detailing the specific implementation. To handle different resolutions of the input point cloud, we divide them into different numbers of patches with a linear scaling (e.g., $M=1024$ input points are divided into $n=64$ point patches), with each patch containing $k=32$ points determined by the KNN algorithm. The PointMamba encoder has $N=12$ vanilla Mamba blocks, each Mamba block featuring $C=384$ hidden dimensions. For the pre-training, we utilize ShapeNetCore chang2015shapenet as the dataset, following previous methods yu2022point ; pang2022masked ; chen2023pointgpt . In addition, we utilize 4 $\times$ Mamba blocks as the decoder to reconstruct the masked point clouds.
Table 1: Object classification on the ScanObjectNN dataset uy2019revisiting . We evaluate PointMamba on three variants, with PB-T50-RS being the most challenging. Overall accuracy (%) is reported. Param. denotes the number of tunable parameters during training. indicates that using simple rotational augmentation dong2022autoencoders for training.
Methods Reference Backbone Param. (M) $\downarrow$ FLOPs (G) $\downarrow$ OBJ-BG $\uparrow$ OBJ-ONLY $\uparrow$ PB-T50-RS $\uparrow$
Supervised Learning Only
PointNet qi2017pointnet CVPR 17 - 3.5 0.5 73.3 79.2 68.0
PointNet++ qi2017pointnet++ NeurIPS 17 - 1.5 1.7 82.3 84.3 77.9
PointCNN li2018pointcnn NeurIPS 18 - 0.6 0.9 86.1 85.5 78.5
DGCNN wang2019dynamic TOG 19 - 1.8 2.4 82.8 86.2 78.1
PRANet ptpranet TIP 21 - - - - - 81.0
MVTN hamdi2021mvtn ICCV 21 - 11.2 43.7 - - 82.8
PointNeXt qian2022pointnext NeurIPS 22 - 1.4 1.6 - - 87.7
PointMLP ma2022rethinking ICLR 22 - 13.2 31.4 - - 85.4
RepSurf-U ran2022surface CVPR 22 - 1.5 0.8 - - 84.3
ADS hong2023attention ICCV 23 - - - - - 87.5
Training from pre-training (Single-Modal)
Point-BERT yu2022point CVPR 22 Transformer 22.1 4.8 87.43 88.12 83.07
MaskPoint liu2022masked CVPR 22 Transformer 22.1 4.8 89.30 88.10 84.30
Point-MAE pang2022masked ECCV 22 Transformer 22.1 4.8 90.02 88.29 85.18
Point-M2AE zhang2022point NeurIPS 22 Transformer 15.3 3.6 91.22 88.81 86.43
PointDif zheng2023point CVPR 24 Transformer - - 93.29 91.91 87.61
Point-MAE+IDPT zha2023instance ICCV 23 Transformer 1.7 7.2 91.22 90.02 84.94
Point-MAE+DAPT zhou2024dynamic CVPR 24 Transformer 1.1 5.0 90.88 90.19 85.08
Point-MAE pang2022masked ECCV 22 Transformer 22.1 4.8 92.77 91.22 89.04
PointGPT-S chen2023pointgpt NeurIPS 23 Transformer 29.2 5.7 93.39 92.43 89.17
PointMamba (ours) - Mamba 12.3 3.1 94.32 92.60 89.31
Training from pre-training (Cross-Modal)
ACT dong2022autoencoders ICLR 23 Transformer 22.1 4.8 93.29 91.91 88.21
Joint-MAE guo2023joint IJCAI 23 Transformer - - 90.94 88.86 86.07
I2P-MAE zhang2023learning CVPR 23 Transformer 15.3 - 94.15 91.57 90.11
ReCon qi2023recon ICML 23 Transformer 43.6 5.3 95.18 93.29 90.63
Figure 5: Classification on ModelNet40 wu20153d . Overall accuracy (%) is reported. The results are obtained from 1024 points without voting.
Methods Param. (M) $\downarrow$ FLOPs (G) $\downarrow$ OA (%) $\uparrow$
Supervised Learning Only
PointNet qi2017pointnet 3.5 0.5 89.2
PointNet++ qi2017pointnet++ 1.5 1.7 90.7
PointCNN li2018pointcnn 0.6 - 92.2
DGCNN phan2018dgcnn 1.8 2.4 92.9
PointNeXt qian2022pointnext 1.4 1.6 92.9
PCT guo2021pct 2.9 2.3 93.2
OctFormer wang2023octformer 3.98 31.3 92.7
with Self-supervised pre-training
Point-BERT yu2022point 22.1 2.3 92.7
MaskPoint liu2022masked 22.1 2.3 92.6
Point-M2AE zhang2022point 12.8 4.6 93.4
Point-MAE pang2022masked 22.1 2.4 93.2
PointGPT-S chen2023pointgpt 29.2 2.9 93.3
ACT dong2022autoencoders 22.1 2.4 93.6
PointMamba (ours) 12.3 1.5 93.6
Figure 6: Few-shot learning on ModelNet40 wu20153d . Overall accuracy (%)$\pm$±the standard deviation (%) without voting is reported.
Methods 5-way 10-way
10-shot 20-shot 10-shot 20-shot
Supervised Learning Only
PointNet qi2017pointnet 52.0$\pm$3.8 57.8$\pm$4.9 46.6$\pm$4.3 35.2$\pm$4.8
PointNet-CrossPoint afham2022crosspoint 90.9$\pm$1.9 93.5$\pm$4.4 84.6$\pm$4.7 90.2$\pm$2.2
DGCNN wang2019dynamic 31.6$\pm$2.8 40.8$\pm$4.6 19.9$\pm$2.1 16.9$\pm$1.5
DGCNN-CrossPoint afham2022crosspoint 92.5$\pm$3.0 94.9$\pm$2.1 83.6$\pm$5.3 87.9$\pm$4.2
with Self-supervised pre-training
Point-BERT yu2022point 94.6$\pm$3.1 96.3$\pm$2.7 91.0$\pm$5.4 92.7$\pm$5.1
MaskPoint liu2022masked 95.0$\pm$3.7 97.2$\pm$1.7 91.4$\pm$4.0 93.4$\pm$3.5
Point-MAE pang2022masked 96.3$\pm$2.5 97.8$\pm$1.8 92.6$\pm$4.1 95.0$\pm$3.0
Point-M2AE zhang2022point 96.8$\pm$1.8 98.3$\pm$1.4 92.3$\pm$4.5 95.0$\pm$3.0
PointGPT-S chen2023pointgpt 96.8$\pm$2.0 98.6$\pm$1.1 92.6$\pm$4.6 95.2$\pm$3.4
ACT dong2022autoencoders 96.8$\pm$2.3 98.0$\pm$1.4 93.3$\pm$4.0 95.6$\pm$2.8
PointMamba (ours) 96.9$\pm$2.0 99.0$\pm$1.1 93.0$\pm$4.4 95.6$\pm$3.2
5.2 Compared with Transformer-based counterparts
This paper aims to unlock the potential of Mamba in point cloud tasks, discussing whether it can be a viable alternative to Transformers. Thus, in the following experiments, we mainly compare with the state-of-the-art vanilla Transformer-based point cloud analysis methods.
Real-world object classification on ScanObjectNN. ScanObjectNN uy2019revisiting is a challenging 3D dataset comprising about 15,000 objects across 15 categories, scanned from real-world indoor scenes with cluttered complexity backgrounds. As shown in Tab. 1, we conduct experiments on three versions of ScanObjectNN (i.e., OBJ-BG, OBJ-ONLY, and PB-T50-RS), each with increasing complexity. When compared with the most convincing Transformer-based method, i.e., Point-MAE liu2022masked , PointMamba surpasses it by 1.55%, 1.38% and 0.27% on OBJ-BG, OBJ-ONLY, and PB-T50-RS respectively while using less computational costs. Besides, we also outperform the SOTA PointGPT-S chen2023pointgpt by 0.93%, 0.17%, 0.14% across three variants on a comparable scale setting. Note that our method follows Occam’s Razor, without auxiliary tasks like generation during fine-tuning used in PointGPT chen2023pointgpt . Furthermore, compared to cross-modal learning methods dong2022autoencoders ; qi2023recon that use additional training data (cross-modal information) or teacher models, which is not a fair comparison, our PointMamba still maintains highly competitive. We mainly want to introduce a new Mamba-based point cloud analysis methods paradigm. Although using some complex designs can bring improvement, they might be heuristics. More importantly, these heuristic designs will decrease the objectivity of the evaluation of our method.
Synthetic object classification on ModelNet40. ModelNet40 wu20153d is a pristine 3D CAD dataset consisting of 12,311 clean samples across 40 categories. As shown in Tab. 6, we report the overall accuracy without adopting the voting strategy. The proposed PointMamba achieves the best results compared with various self-supervised Transformer-based methods liu2022masked ; yu2022point ; chen2023pointgpt . In particular, PointMamba surpasses Point-MAE liu2022masked and PointGPT-S chen2023pointgpt by 0.4% and 0.3% respectively. It is worth noting that the single-modal-learned PointMamba achieves comparable results with the cross-modal-based ACT dong2022autoencoders while significantly reducing parameters and FLOPs about 44% and 38%, respectively. Additionally, PointMamba demonstrates competitive performance against elaborately designed Transformer models like OctFormer wang2023octformer .
Table 2: Part segmentation on the ShapeNetPart yi2016scalable . The mIoU for all classes (Cls.) and for all instances (Inst.) are reported.
Methods Cls. mIoU (%) $\uparrow$ Inst. mIoU (%) $\uparrow$
Supervised Learing Only
PointNet qi2017pointnet 80.39 83.7
PointNet++ qi2017pointnet++ 81.85 85.1
DGCNN wang2019dynamic 82.33 85.2
APES wu2023attention 83.67 85.8
with Self-supervised pre-training
Transformer yu2022point 83.4 85.1
OcCo yu2022point 83.4 85.1
MaskPoint liu2022masked 84.6 86.0
Point-BERT yu2022point 84.1 85.6
Point-MAE pang2022masked 84.2 86.1
PointGPT-S chen2023pointgpt 84.1 86.2
ACT dong2022autoencoders 84.7 86.1
PointMamba (ours) 84.4 86.2
Few-shot learning. We further conduct few-shot experiments on ModelNet40 wu20153d to demonstrate our few-shot transfer ability. Consistent with prior studies yu2022point , we utilize the "$n$-way, $m$-shot" setup, where $n\in\{5,10\}$ denotes the category count and $m\in\{10,20\}$ represents the samples per category. Following standard procedure, we carry out 10 separate experiments for each setting and reported mean accuracy along with the standard deviation. As indicated in Tab. 6, our PointMamba shows competitive results with limited data, e.g., +1.0% mean accuracy compared to the cross-modal method ACT dong2022autoencoders on the 5-way 20-shot split.
Part segmentation on ShapeNetPart. Part segmentation on ShapeNetPart yi2016scalable is a challenging task that aims to predict a more detailed label for each point within a sample. As shown in Tab. 2, we report mean IoU (mIoU) for all classes (Cls.) and all instances (Inst.). Our PointMamba model demonstrates highly competitive performance compared to the Transformer-based counterparts liu2022masked ; chen2023pointgpt ; yu2022point . These impressive results further prove the potential of SSM in the point cloud analysis tasks.
Figure 7: Different variant of PointMamba. (a) Directly removing the SSM part. (b) Replacing SSM with attention. (c) Replacing SSM with MLP. (d) Ours PointMamba with Selective SSM.
Figure 8: The effect of each component.
Hilbert Trans-Hilbert Order indicator OBJ-BG OBJ-ONLY
Random - 92.60 90.18
- - 92.94 91.05
- - 93.46 91.74
- 93.80 91.91
94.32 92.60
Figure 9: The effect of different scanning curves.
Scanning curve OBJ-BG OBJ-ONLY
Random 92.60 90.18
Hilbert and Trans-Hilbert 94.32 92.60
Z-order and Trans-Z-zorder 93.29 90.36
Hilbert and Z-order 93.29 90.88
Trans-Hilbert and Trans-Z-order 93.29 91.91
Figure 10: The effect of Selective SSM.
Setting Param. OBJ-BG OBJ-ONLY
w/ Identity 11.4 93.80 91.57
w/ Attention 39.8 92.77 91.22
w/ MLP 18.5 93.29 91.22
w/ Selective SSM 12.3 94.32 92.60
Figure 11: The effect of Order indicator.
Setting OBJ-BG OBJ-ONLY
None 93.80 91.91
Using the same indicator 93.29 90.19
Using different indicator 94.32 92.60
5.3 Analysis and ablation study
To investigate the architecture design, we conduct ablation studies on ScanObjectNN uy2019revisiting with both pre-training and fine-tuning. Default settings are marked in gray.
The structural efficiency. We first discuss the efficiency of our method. To fully explore the potential of processing the long point tokens (sequence), we gradually increase the sequence length until the GPU (NVIDIA A800 80GB) memory explodes. The comprehensive efficiency comparisons are present in Fig. 1(b)-(d), where Compared with the most convincing Transformer-based method liu2022masked , our PointMamba demonstrates significantly improved inference speed and reduce the GPU usage and FLOPs, especially when facing the long sequence. For example, when the length increases to more than 32,768, we outperform PointMAE by 30.2$\times$, 24.9$\times$, and 5.2$\times$ in terms of inference speed, GPU memory, and FLOPs, respectively. More importantly, even presenting impressive efficiency, we still achieve impressive performance on various point cloud analysis datasets.
The effect of each component. We then study the effectiveness of the proposed components of PointMamba as shown in Tab. 9. We can make the following observations: 1) Directly utilizing random serialization, PointMamba only achieves 92.26% and 90.18% overall accuracy on OBJ-BG and OBJ-ONLY, respectively. It is reasonable as Mamba is hard to model the unstructured point clouds due to its unidirectional modeling. 2) By introducing the locality-preserved Hilbert or Trans-Hilbert scanning, PointMamba’s ability to capture sequence information is enhanced, leading to performance improvements compared to random serialization. Further applying both Hilbert and Trans-Hilbert scanning curves, PointMamba surpasses the random serialization by 1.20% and 1.73% on two datasets, respectively. 3) By using the order indicator to maintain the distinct characteristics of the two different scanning strategies, we achieve notable improvement, resulting in 94.32% and 92.60% on OBJ-BG and OBJ-ONLY, respectively. Note that the order indicator is extremely light (only 1.5k parameters), which will not introduce additional computational costs.
The effect of different scanning curves. We further explore the effect of using different scanning curves to construct serialized point tokens. Specifically, we select two widely used space-filling curves, including Hilbert and Z-order, along with their transposed variants, i.e., Trans-Hilbert and Trans-Z-order. As listed in Tab. 9, we empirically find that serializing point clouds with space-filling curves scanning can achieve better performance compared to random sequences. We argue that scanning sequences along a specific pattern of spatial locations offers a more logical sequence modeling order for SSM. We choose the combination of Hilbert and Trans-Hilbert for PointMamba due to their superior locality-preserving properties.
The effect of Selective SSM. The key of S6 models or Mamba gu2023mamba is the SSM with the selective mechanism. We prove that as a unidirectional modeling method, SSM can be analogous to masked self-attention, ensuring each position can only attend to previous positions (the detailed proof can be found in the Appendix A.1). Thus, as shown in Tab. 11, we analyze the effect of selective SSM by removing it (i.e., identity setting) or replace with masked self-attention or MLP (an illustration is shown in Fig. 7). Compared with the identity setting, the selective SSM brings notable improvement, indicating the effectiveness of introducing global modeling from SSM. Note that while a very recent method, MambaOut yu2024mambaout , thinks the SSM of Mamba might negatively impact image classification tasks, our findings demonstrate that this is not the case for point cloud analysis tasks. Another interesting thing is that when Selective SSM is replaced with masked self-attention, the performance is even lower than that of the identity setting. We argue the main reason is that masked self-attention is hard to combine with the Gated MLP shazeer2020glu used in default Mamba, leading to optimized difficulty, which might need to be explored in the future.
Analysis on order indicator. This part analyzes the effect of the order indicator. PointMamba applies Hilbert and Trans-Hilbert to recognize the key points, obtaining two types of serialized point tokens $\bm{E}_{0}^{h}$ and $\bm{E}_{0}^{h^{\prime}}$. The order indicator is used to indicate the scanning strategy. As shown in Tab. 11, using two different order indicators can improve 1.20% and 1.03% compared to no indicator on OBJ-BG and OBJ-ONLY, respectively. However, using the same order indicator for both types of sequences without distinguishing between different scanning strategies does not yield positive results.
5.4 Limitation
Although PointMamba achieves promising results, there are some limitations: 1) We only focus on the point cloud analysis task in this paper while designing a unified Mamba-based foundation model for various 3D vision tasks (e.g., 3D object classification/detection/segmentation) is a more appealing direction. 2) We only use the point clouds as training data while combining them with 2D images or language knowledge to improve the performance, which is also worthy of exploration. We left these in our future work.
6 Conclusion
In this paper, we present an elegant, simple Mamba-based method named PointMamba for point cloud analysis. PointMamba utilizes a space-filling curve-based point tokenizer and a plain, non-hierarchical Mamba architecture to achieve global modeling with linear complexity. Despite its structural simplicity, PointMamba delivers state-of-the-art performance across various datasets, significantly reducing computational costs in terms of GPU memory and FLOPs. PointMamba success highlights the potential of SSMs, particularly Mamba, in handling the complexities of point cloud data. As a newcomer to point cloud analysis, PointMamba is a promising option for constructing 3D vision foundation models, and we hope it can offer a new perspective for the field.
Appendix A Theoretically Analysis
A.1 Closer look at Selective SSM
As described in Sec. 3, Selective SSM gu2023mamba considers parameters $\bm{B},\bm{C},\bm{\Delta}$ in Eq. 2 as functions of the input. To be specific, given the input sequence $\hat{\bm{x}}=\left[x_{1},\cdots,x_{t},\cdots,x_{L}\right]\in\mathbb{R}^{L\times C}$, the per-time matrices $\bm{B}_{t},\bm{C}_{t},\bm{\Delta}_{t}$ can be computed as follows:
$\bm{B}_{t}=L_{B}(x_{t}),\quad\bm{C}_{t}=L_{C}(x_{t}),\quad\bm{\Delta}_{t}=\mathrm{softplus}\left(L_{\Delta}(x_{t})\right),$ (5)
where $L_{B},L_{C},L_{\Delta}$ are linear projection layers, and $\mathrm{softplus}(x)=\log(1+e^{x})$. The matrices $\overline{\bm{A}}_{t},\overline{\bm{B}}_{t},\bm{C}_{t},\bm{\Delta}_{t}$ can be obtained by taking Eq. 5 into Eq. 2.
To simplify, we ignore the residual connection $\bm{D}$ and expand Eq. 1, the output $\hat{\bm{y}}=\left[y_{1},\cdots,y_{t},\cdots,y_{L}\right]\in\mathbb{R}^{L\times C}$ can be computed below:
$y_{t}=\bm{C}_{t}h_{t},\quad h_{t}=\sum_{i=1}^{t}\left(\prod_{j=i+1}^{t}\overline{\bm{A}}_{j}\right)\overline{\bm{B}}_{i}x_{i},$ (6)
which can be further described in matrix form below:
$\begin{bmatrix}h_{1}\\ h_{2}\\ \vdots\\ h_{t}\\ \end{bmatrix}=\begin{bmatrix}\overline{\bm{B}}_{1}&0&\cdots&0\\ \overline{\bm{A}}_{2}\overline{\bm{B}}_{1}&\overline{\bm{B}}_{2}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ {\textstyle\prod_{i=2}^{t}}\overline{\bm{A}}_{i}\overline{\bm{B}}_{1}&{\textstyle\prod_{i=3}^{t}}\overline{\bm{A}}_{i}\overline{\bm{B}}_{2}&\cdots&\overline{\bm{B}}_{t}\end{bmatrix}\begin{bmatrix}x_{1}\\ x_{2}\\ \vdots\\ x_{t}\\ \end{bmatrix}.$ (7)
For an intuitive understanding, Eq. 7 resembles the self-attention mechanism with a mask $\bm{M}$, specifically causal self-attention. In this context, $\bm{M}$ is a lower triangular matrix with elements set to 1. To further exam this, consider the transfer matrix $\bm{W}$ between $\hat{y}$ and $\hat{x}$, i.e., $\left(\hat{\bm{y}}=\bm{W}\hat{\bm{x}}\right)$:
$\bm{W}_{i,j}$ $=\bm{C}_{i}\left(\prod_{k=j+1}^{i}\overline{\bm{A}}_{k}\right)\overline{\bm{B}}_{j}$ (8)
$=\bm{C}_{i}\left(\prod_{k=j+1}^{i}\exp\left(\bm{\Delta_{k}A}\right)\right)\overline{\bm{B}}_{j}$ (9)
$=\bm{C}_{i}\exp\left(\sum_{k=j+1}^{i}\bm{\Delta_{k}A}\right)\overline{\bm{B}}_{j}$ (10)
$\approx\bm{C}_{i}\exp\left(\mathop{\sum}_{k=j+1\atop L_{\Delta}\left(x_{k}\right)>0}^{i}L_{\Delta}\left(x_{k}\right)\bm{A}\right)\overline{\bm{B}}_{j},$ (11)
where $\bm{W}_{i,j}$ represents the element in the $i$-th row and $j$-th column, the approximation in Eq. 11 is done using $\mathrm{ReLU}$ instead of $\mathrm{softplus}$. Consider the notation below:
$\bm{Q}_{i}:=\bm{C}_{i},\qquad\bm{T}_{i,j}=\exp\left(\mathop{\sum}_{k=j+1\atop L_{\Delta}\left(x_{k}\right)>0}^{i}\left(L_{\Delta}\left(x_{k}\right)\bm{A}\right)\right),\qquad\bm{K}_{j}=\left(\overline{\bm{B}}_{j}\right)^{T}.$ (12)
Thus, the Eq. 11 can be simplified to:
$\bm{W}_{i,j}$ $\approx\bm{Q}_{i}\bm{T}_{i,j}\bm{K}_{j}^{T}.$ (13)
This shows that the Selective SSM captures the influence of $x_{i}$ and $x_{j}$ using $\bm{Q}_{i}$ and $\bm{K}_{j}$, respectively, while $\bm{T}_{i,j}$ molding the token significance from $x_{i}$ to $x_{j}$. Note that $i\leq j$ because $\bm{W}$ is a lower triangular matrix, indicating a strong relationship with causal self-attention brown2020language ; ali2024hidden .
A.2 Global modeling of PointMamba
In this subsection, we explain the global modeling of our PointMamba. Let’s consider the total input sequence as $\left[\hat{\bm{l}}_{1};\hat{\bm{l}}_{2}\right]=\left[x_{1},,\cdots,x_{l/2};x_{l/2+1},\cdots,x_{l}\right]$ where the sequence has a length $l$ and $l$ is even. $\hat{\bm{l}}_{1}$ comes from Hilbert serialization and the other half, $\hat{\bm{l}}_{2}$, comes from Trans-Hilbert. The large matrix in Eq. 7 can be represented as a partitioned matrix $\begin{bmatrix}\bm{X}&\bm{0}\\ \bm{Y}&\bm{Z}\end{bmatrix}$ as below:
[…]
Note that the block $\bm{Y}$, highlighted in gray, is associated with both $\hat{\bm{l}}_{1}$ (from $\overline{\bm{B}}$) and $\hat{\bm{l}}_{2}$ (from $\overline{\bm{A}}$), denoted as $\bm{Y}(\hat{\bm{l}}_{1},\hat{\bm{l}}_{2})$. The blocks $\bm{X}$ and $\bm{Z}$ only relate to half of the sequence, denoted as $\bm{X}(\hat{\bm{l}}_{1})$ and $\bm{Z}(\hat{\bm{l}}_{2})$, respectively. Thus, the hidden space output $\left[\hat{\bm{h}}_{1};\hat{\bm{h}}_{2}\right]^{T}=\left[h_{1},\cdots,h_{l/2};h_{l/2+1},\cdots,h_{l}\right]^{T}$ can be compressed as below:
$\begin{bmatrix}\hat{\bm{h}}_{1}\\ \hat{\bm{h}}_{2}\end{bmatrix}=\begin{bmatrix}\bm{X}(\hat{\bm{l}}_{1})\hat{\bm{l}}_{1}\\ \bm{Y}(\hat{\bm{l}}_{1},\hat{\bm{l}}_{2})\hat{\bm{l}}_{2}+\bm{Z}(\hat{\bm{l}}_{2})\hat{\bm{l}}_{2}\end{bmatrix}.$ (15)
As in Fig. 3 and Eq. 15, the serialized points from Trans-Hilbert can receive global information from Hilbert serialization.
Appendix B Concurrent Related Works
Table 3: Classification performance comparisons with other state space model methods on three variants of the ScanObjectNN uy2019revisiting . All results are reported without voting.
Method Param. (M) FLOPs (G) ScanObjectNN
OBJ_BG OBJ_ONLY PB_T50_RS
Point Cloud Mamba zhang2024point 34.2 45.0 - - 88.10
Mamba3D han2024mamba3d 16.9 3.9 93.12 92.08 88.20
PoinTramba wang2024pointramba 19.5 - 92.30 91.30 89.10
PointMamba (ours) 12.3 3.1 94.32 92.60 89.31
Some works on state space models for point cloud analysis appeared recently. This section discusses the differences between these methods and our PointMamba.
Point Cloud Mamba (PCM) zhang2024point combines an improved Mamba module, i.e., Vision Mamba zhu2024vision , with PointMLP ma2022rethinking (a strong point cloud analysis method), and incorporates consistent traverse serialization at each stage. To enhance Mamba’s capability in managing point sequences with varying orders, PCM introduces point prompts that convey the sequence’s arrangement rules. While these techniques improve the performance of state space models, they also introduce additional computational overhead and complexity in design.
Mamba3D han2024mamba3d is another recently proposed method. To obtain better global features, Mamba3D introduces an enhanced Vision Mamba zhu2024vision block, which includes both a token forward SSM and a backward SSM that operates on the feature channel. It proposes a Local Norm Pooling block to extract local geometric features.
PointTramba wang2024pointramba introduces a hybrid approach that integrates Transformers and Mamba. It segments point clouds into groups and utilizes Transformers to capture intra-group dependencies, while Mamba models inter-group relationships using a bi-directional, importance-aware ordering strategy.
As shown in Tab. 3, our PointMamba surpasses these concurrent methods by offering superior performance with reduced computational overhead. Note that our method is extremely simple and without complex design, which utilizes the vanilla Mamba block and abstains from incorporating modular designs from other baselines, thereby maintaining simplicity and efficiency in our approach. We believe such a method can better illustrate the potential of SSM in point cloud analysis tasks.
Appendix C More experimental resutls
C.1 Implement details
Table 4: Implementation details for pre-training and downstream tasks.
Configuration Pre-training Classification Segmentation
ShapeNetCore ModelNet40 ScanObjectNN ShapeNetPart
Optimizer AdamW AdamW AdamW AdamW
Learning rate 1e-3 3e-4 5e-4 2e-4
Weight decay 5e-2 5e-2 5e-2 5e-2
Learning rate scheduler cosine cosine cosine cosine
Training epochs 300 300 300 300
Warmup epochs 10 10 10 10
Batch size 128 32 32 16
Num. of encoder layers $N$ 12 12 12 12
Num. of decoder layers 4 - - -
Input points $M$ 1024 1024 2048 2048
Num. of patches $n$ 64 64 128 128
Patch size $k$ 32 32 32 32
Augmentation Scale&Trans Scale&Trans Rotation -
Pre-training Details. The ShapeNetCore dataset chang2015shapenet is used for pre-training, including $\sim$51K clean 3D sample across 55 categories. The 1,024 input points are divided into 64 point patches, with each patch consisting of 32 points. The pre-training process includes 300 epochs, with a batch size of 128. More detail can be found in Tab. 4.
Figure 12: The details of our PointMamba for segmentation task.
Downstream tasks Details. Fig. 2 shows the pipeline of PointMamba for classification tasks. We report the overall accuracy without voting on the challenging ScanObjectNN uy2019revisiting using 2,048 input points, and on ModelNet40 wu20153d using 1,024 input points. For segmentation on ShapeNetPart yi2016scalable , as shown in Fig. 12, we use random, Hilbert, and Trans-Hilbert serializations, with order indicators applied on Hilbert/Trans-Hilbert serializations. Features from the 3-rd, 7-th, and last layer are pooled as global features after a simple feature fusion. These global features are then concatenated with per-point features and sent to the segmentation head. More detail can be found in Tab. 4.
C.2 Additional ablation study
Figure 13: The effect of masking strategy. The pre-training loss (× 1000) along with fine-tuning accuracy (%) are reported.
Masking ratio Loss OBJ-BG OBJ-ONLY
0.4 2.01 92.60 90.70
0.6 1.97 94.32 92.60
0.8 2.33 93.46 90.17
0.9 2.00 92.43 91.05
Figure 14: The effect of classification token. Fine-tuning accuracy (%) are reported.
Methods OBJ-BG OBJ-ONLY
Before the sequence 93.63 91.05
After the sequence 94.32 90.19
Middle the sequence 93.98 90.71
Max Pool 93.39 91.36
Average Pool 94.32 92.60
In this section, we do additional ablation studies on several hyper-parameters.
Masking strategy for pre-training. By employing a serialization-based mask modeling paradigm, our PointMamba achieves superior performance. To find a proper masking strategy for our method, we compare two types of masking with varying ratios. The block masking yu2022point masks geometrically proximate point cloud patches, leading to a more challenging reconstruction target. In Tab. 14, we experimentally find that masking 60% of point patches by randomly choosing can achieve good performance.
Usage of classification token. Previous works dosovitskiy2020image ; yu2022point ; pang2022masked often use a classification token $[\mathtt{CLS}]$ as a global token for classification. As in Tab. 14, we find that without $[\mathtt{CLS}]$ and utilizing only the average pooling of the final block’s output yields the best results for PointMamba.
Appendix D Qualitative Analysis
D.1 Mask modeling visualization
As in Sec. 4.2, we customize a simple yet effective serialization-based mask modeling paradigm. By randomly masking about 60% of serialization-based point tokens, an asymmetric vanilla Mamba autoencoder is utilized to extract the point feature, with a simple prediction head for reconstruction. In Fig. 15, we present qualitative results of mask modeling on ShapeNet validation set. Despite a 60% masking ratio, our PointMamba effectively reconstructs the masked patches, providing a strong self-supervised knowledge for downstream tasks.
Figure 15: The qualitative results of mask predictions of our PointMamba on ShapeNet validation set.
Figure 16: The qualitative results of part segmentation of our PointMamba on ShapeNetPart.
D.2 Part segmentation visualization
In this subsection, we present the qualitative results for part segmentation on the ShapeNetPart validation set, including both the ground truth and the predicted results. As in Fig. 16, our PointMamba shows highly competitive results on part segmentation.